OpenStudy (anonymous):

Is 7/12 greater or less than or equal to 2/3? Help, Thank you if you helped

OpenStudy (mrnood):

in order to be able to compare the 2 numbers easily they need to have the same denominator. The easiest way to find a common denominator is to multiply the 2 denominators together (this in not necessarily the LOWEST common denominator - but it doesn't matter in this case). SO what is the denominator if you multiply the two denominators of 7/12 and 2/3 ?

OpenStudy (anonymous):

36? Or does the 2/3 equal 8/12?

OpenStudy (mrnood):

36 is correct so: \[\frac{ 7 }{ 12}= \frac{ ? }{ 36 }\] \[\frac{ 2 }{ 3 } = \frac{ ? }{ 36}\] You can then easily see which number is bigger

OpenStudy (mrnood):

(and BTW - you are correct that 2/3 = 8/12. That is the quickest and most direct solution. But the way I suggest works for all combinations of denominator)

OpenStudy (anonymous):

Do I multiply 12 x 7 and 12 x 12, then 12 x 2 and 12 x 3(36) and this will show me the greater fraction?

OpenStudy (mrnood):

no to convert 12 to 36 you multiply by 3 - so multiply 7 by 3 to convert 3 to 36 you multiply by 12 - so multiply 2 by 12...